2. Taxes paid for a gliven income level to do her taxes. She is single and lives in San Francisco. Beth earned $110,000 in taxable income in 2015. She reviews the following table, which shows the IRS tax rates for a single taxpayer in 2015 The Tax Rate Is... (Percent) 10.0 15.0 25.0 28.0 3.0 35.0 39.6 On Annual Taxable Income... Up to $9,225 From $9,225 80 537,450 From $37,450 to $90,750 From $90,750 to $189,300 From $189,300 to $411,500 From $411,500 to $413,200 Over $413,200 Based on the IS table, Beth calculates that her marginal tax rate is when her annual taxable income is $110.000 Beth caloulates that she owes n income taxes for 20 ,based on the annual income level and the amount of taxes she owes o fter iguring out whalt she owes in taxes in 2015, Beth deides to ask an accountant for tax advice. The accountant claims that he has found a lega way to shelter $2,000 of taxable income from the redrai government. The maximum amount thar Beth is willieg to pay to learn this strategy and reduce her taxable income by $2.000
Answer #1

Marginal Tax rate = 28%.

The marginal tax rate is the tax paid on the last dollar earned. Since her income is 110,000 the last tax she would face is 28%. Thus the last dollar will be taxed at 28%.

She owes $23871.25 in taxes

Up to 9225 - 10% tax = 922.5

9225 - 37450 - 15% tax = 37450-9225 at 15% = 4233.75

37450 - 90750 - 25% tax = 90750-37450 at 25% = 13325

90750 - 189300 - 28% tax = 110,000-90750 at 28% = 5390

Total Tax = 23871.25

Average Tax rate = 23871.25/110,000 = 21.7%

The accountant can save her 2000 in taxes, thus maximum she would be willing to pay would be $2000.